OpenStudy (anonymous):

Which of the following is a polynomial with roots 5, −6, and 7? A. f(x) = x3 − 6x2 − 30x + 37 B. f(x) = x3 − 30x2 − 37x + 42 C. f(x) = x3 − 6x2 − 37x + 210 D. f(x) = x3 − 30x2 − 42x + 210

OpenStudy (vjt):

I don't know the answer and I forgot how to do these, but someone on yahoo had a similar question with different numbers. Maybe the answer they got can help you. https://answers.yahoo.com/question/index?qid=20151029103535AA5FiWd If you consider the expansion of (x-4)(x-6)(x+7), you see that the constant term must be .. (-4)(-6)(+7) = 168 This rules out choices A and C. If you consider the coefficient of the x^2 term, you realize it must be (-4) + (-6) + (7) = -3. This rules out choice D and suggests choice B. _____ If you choose to evaluate the polynomial at each of the roots, you should do it as .. x^2 + bx^2 + cx + d .. = ((x + b)x + c)x + d ... "Horner's form" of the polynomial Evaluation in this form is similar to synthetic division. For the first one, using x=4, you get .. ((4 -3)4 - 24)4 + 42 = (4-24)4 + 42 = -20*4+42 (obviously not zero)